A piece of wire 44cm long is cut into two parts and each part is bent to form a square. If the total area of the two squares is 65 square centimeter, find the perimeter of the two squares.
Answer by lwsshak3(11628)

A piece of wire 44cm long is cut into two parts and each part is bent to form a square. If the total area of the two squares is 65 square centimeter, find the perimeter of the two squares.
...
let x=side of one square in cm
perimeter of one square=4x
let y=side of other square in cm
perimeter of other square=4y
..
x^2+y^2=65
..
4x+4y=44
x+y=11
y=11-x
..
x^2+(11-x)^2=65
x^2+121-22x+x^2=65
2x^2-22x+56=0
x^2-11x+28=0
(x-4)(x-7)=0
x=4 or x=7
y=11-4=7 or y=11-7=4
ans:
Perimeter of one square= 4*4=16 cm
Perimeter of other square = 4*7=28 cm
If the perimeter of one square is 16 cm, perimeter of the other square would be 28 cm or vice-versa